Weaver Company’s predetermined overhead rate is $15.00 per direct labour-hour, and its direct labour wage rate is $10.00 per hour. The following information pertains to Job A-200:

    
  Direct materials $315
  Direct labour $210

Required:

1. What is the total manufacturing cost assigned to Job A-200?

2. If Job A-200 consists of 40 units, what is the average cost assigned to each unit included in the job?

Score answer Visit question map

Question 1 of 2 Total
Answer #1

1.

Direct labor hours used in Job A-200= Direct labor cost/ Direct labor rate

= 210/10

= 21

Overhead applied = Direct labor used x Predetermined overhead rate

= 21 x 15

= $315

Job Cost Sheet ( Job A-200)

Direct material 315
Direct labor 210
Applied overhead 315
Total Cost $840

2.

Cost per unit = Total cost/ Number of units

= 840/40

= $21

Kindly comment if you need further assistance. Thanks‼!